OpenStudy (anonymous):

Tell whether the equations are parallel, perpendicular, or neither.y = -1/4x + 10 and -2x + 8y = 6

OpenStudy (anonymous):

parallel.

OpenStudy (anonymous):

because they have the same slope.

OpenStudy (anonymous):

@kumarsajan What are the slopes?

OpenStudy (anonymous):

y = -1/4x + 10 ----> slope here is -1/4 -2x + 8y = 6 --> you must convert it to slope intercept form ( y = mx + b) so divide the second equation by 8. (-2x +8y =6)/8 -1/4x + y = 6/8 y = 1/4x + 6/8 aaaaaand i was wrong the answer is neither. :)))